OpenStudy (anonymous):

A boy is 6 ft. tall. The distance from the boy to a mirror is 8 ft. From the mirror to the house is 16 ft. How high is the top of the house? x = ft.

OpenStudy (amistre64):

sounds like there is a picture involved?

OpenStudy (amistre64):

otherwise these dont really correspond to anything

OpenStudy (anonymous):

or we need just a little more info..

OpenStudy (anonymous):

there is no picture but let me see

OpenStudy (anonymous):

all it say is use similiar triangles there is no picture

OpenStudy (amistre64):

cant reallly tell what is sposed to be similar; there is no reference to it in the question itself

OpenStudy (amistre64):

i think I see it tho there is a mirror onthe ground; and the boy is looking into it and seeing the top of the house...

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then similar tris gives you a house at 12 feet

OpenStudy (amistre64):

might be wrong in that

OpenStudy (anonymous):

yep your right :)

OpenStudy (amistre64):

6/8 = h/16 6(16)/8 = h

OpenStudy (amistre64):

ok :) had me doubts cause i tend to do em in me head lol